# LaDTalk: Latent Denoising for Synthesizing Talking Head Videos with High Frequency Details

### v2

@@ Abstract
Identity-specific audio-driven talking head generation (THG) is crucial for applications in filmmaking and virtual reality. Despite significant progress in existing end-to-end methods, they often struggle to generate videos with high-frequency details due to limited expressivity. In this paper, we tackle this challenge by introducing LaDTalk, a framework designed to generate photorealistic talking head videos. Specifically, we redefine the task of ID-specific THG as temporally consistent face deblurring, aiming to restore audio-synchronized but blurred faces from one-shot base models. This approach allows LaDTalk to leverage the strengths of both one-shot methods and ID-specific approaches, enabling the synthesis of lip-synchronized talking head videos with high fidelity and detailed textures. To achieve this, we theoretically establish the noise tolerance of Vector Quantised Auto Encoders (VQAEs) using Lipschitz Continuity theory. Building on this analysis, we introduce a plug-and-play Space-Optimized VQAE (SOVQAE) that recovers fine-grained textures while ensuring temporal consistency. Extensive experiments demonstrate that LaDTalk achieves state-of-the-art performance in video quality and out-of-domain lip synchronization accuracy when integrated with Wav2Lip as the base model. Furthermore, SOVQAE enhances the performance of various base models, confirming its versatility and effectiveness.
@@ 1 Introduction
The generation of photo-realistic, speech-driven ID-specific talking head videos has significant potential across various domains, including filmmaking [41], virtual reality [42], and digital avatar creation [43]. The goal of this work is to synthesize high-fidelity talking head videos that achieve precise lip synchronization while preserving fine-grained details such as hair, facial wrinkles, moles, eyelashes, and the intricate contours of the lips. These high-frequency details are critical for enhancing realism in synthesized videos.
One-shot techniques [1, 12, 11, 44, 51] have made significant progress in lip synchronization but often struggle to maintain temporal consistency and visual fidelity, leading to issues such as tooth flickering and inconsistent lip thickness. These inconsistencies arise from multi-ID training paradigms that rely heavily on large-scale datasets [10, 56].
To address these challenges, NeRF-based approaches [2, 3, 39, 14, 52] have been introduced, demonstrating impressive results in preserving high-fidelity identities. However, these methods, which are primarily based on Multilayer Perceptrons (MLPs), encounter difficulties in capturing high-frequency details due to the phenomenon known as “spectral bias” [5, 6, 59]. Recent studies on the neural tangent kernel (NTK) [57, 4] reveal that spectral bias arises from the pathological distribution of NTK’s eigenvalues, where most eigenvalues are very small, limiting the convergence speed for high-frequency components.
In contrast, our work focuses on addressing the challenge of generating high-frequency details in ID-specific talking head videos. Specifically, we introduce LaDTalk, an effective framework designed to enhance the output quality of Wav2Lip [1]. By treating its outputs as blurred approximations of ideal results, we redefine the task as one of temporally consistent face deblurring. This design leverages the robust audio-lip synchronization capabilities of one-shot models while incorporating advanced restoration techniques for high-frequency details.
A key challenge in this approach lies in ensuring temporal consistency during denoising while preserving high-fidelity details. To address this challenge, we analyze the forward dynamic process of Vector Quantised Auto Encoders (VQAEs) [^1] We use the term VQAE to distinguish it from VQGAN [21], which combines a VQAE with a transformer network and emphasizes image synthesis. Here, we focus on the auto-encoder nature of VQAE. in the latent space using the Lipschitz Continuity theory of neural networks [9, 8]. Based on this analysis, we theoretically prove the noise robustness of VQAEs (see Sec. 3). This novel theory demonstrates that VQAEs can preserve high-frequency detail information within a discrete codebook while exhibiting noise tolerance, as illustrated in Fig. 2. Building on this insight, we propose a codebook regularization loss to enhance the noise robustness of VQAEs (termed Space-Optimized VQAE, SOVQAE) and facilitate temporally consistent face deblurring.
Figure 1: (a) Illustrates Voronoi regions in 2D and explains the denoising mechanism in VQ space. Given a mapping point $P$, there exists an open disk $\{{\bf x}\in V_{A}\cup V_{B}\cup V_{C}:\|{\bf x}-P\|_{2}<s\}$, where all points in the disk share the same anchor point $A$. (b) Demonstrates the robustness of VQAE under slight input disturbances. When image perturbations occur, the VQ mechanism correctly matches the appropriate codebook vector in the latent space, provided the affected vector remains within the correct Voronoi region (e.g., the yellow-red vector in the figure).
More importantly, SOVQAE is a plug-and-play model that can be integrated into various base models to enhance their performance. Our experiments show that integrating SOVQAE improves the visual quality metrics across different ID-specific base models, confirming its versatility and effectiveness. Especially, when integrated with Wav2Lip [1], LaDTalk surpasses state-of-the-art (SOTA) methods in terms of lip synchronization accuracy, video clarity, and high-frequency detail preservation. Comprehensive evaluations highlight LaDTalk’s superiority in multiple metrics, making it a promising solution for high-fidelity talking head generation.
In summary, this work makes the following contributions:
  • To our knowledge, this study pioneers the exploration of noise robustness in VQAE through the lens of Lipschitz Continuity theory.
  • We propose SOVQAE, a simple and effective plug-and-play model for ID-specific talking head generation that enhances high-fidelity and high-frequency details, supported by rigorous theoretical analysis.
  • Through extensive experiments, we demonstrate the validity and superiority of LaDTalk over SOTA methods across multiple metrics.
@@ 2 Related Work
2.1 Audio-Driven Talking Heads Generation
One-shot Methods, as delineated in [1, 11, 12, 44, 51, 15, 53, 54], harness a plethora of talking video clips to capture the generalizability across multiple faces, thereby attaining commendable lip-synchronization performance. For instance, Wav2Lip [1] pioneers the integration of a lip-sync expert to synchronize audio segments with generated lip movements. DINet [12], in accordance with audio features, induces spatial deformations on reference image feature maps, facilitating few-shot generation capabilities. Despite these advancements, these methods continue to grapple with maintaining identity consistency across video frames. To counter this limitation, ID-specific NeRF-based methods [2, 39, 14, 3, 52, 55] are proposed, focusing exclusively on identity-specific data for training. Although this approach mitigates issues of identity instability, the constrained volume of training data hampers their cross-audio synchronization capabilities. To this end, GeneFace [14] and SyncTalk [2] enhance their audio generalization performance by leveraging pre-trained audio-to-landmarks predictors and audio encoders [1] respectively. Although these initiatives successfully bolster the cross-audio generalization of NeRF methods, instances of cross-audio desynchronization persist. Building upon the foundational capabilities of Wav2Lip within our pipeline, LaDTalk achieves superior cross-audio synchronization performance characteristic of one-to-one methods. Moreover, our approach also excels in visual quality, attributable to the temporally consistent latent denoising prowess of our SOVQAE.
2.2 Face Restoration
Blind Face Restoration [30, 20, 31, 32, 33, 34, 35, 36, 37] endeavors to rectify diverse degradations in facial imagery captured in non-ideal conditions, including compression, blur, and noise. Given the unknown nature of the specific degradations, this task is inherently ill-posed. The prevalent approach to address this challenge is to integrate prior knowledge within an encoder-decoder framework: GFPGAN [36] utilizes a pre-trained face GAN’s rich priors for blind face restoration. Yang et al. [37] integrate a GAN into a U-shaped DNN for enhanced face restoration. Codeformer [36] uses a learned codebook to predict codes, reducing restoration uncertainty. In our approach, we harness a well-trained VQAE as a facial prior to enhance high-frequency restoration from low-quality face images. Through rigorous theoretical analysis and empirical experiments, we establish that the SOVQAE is an effective plug-and-play model for achieving identity-specific, temporally-consistent face restoration, thereby enabling the generation of high-quality talking faces.
2.3 Vector Quantization (VQ) Family
Initially, VQ emerged as a technique within the Signal Processing community [16, 17, 18, 19]. Mathematically, VQ maps signal spaces $\mathbb{R}^{K}$ onto a finite set of Voronoi regions $\{{V_{n}}\}_{n=1}^{N}$, each associated with an anchor vector $\{{\bf c}_{n}\}_{n=1}^{N}\subset\mathbb{R}^{c}$:
@@ 2.3 Vector Quantization (VQ) Family
where $\|\cdot\|_{2}$ denotes the Euclidean distance. Recently, Oord et al. [7] integrated the VQ operation into generative models, sparking a trend of discrete latent representation in the field [20, 15, 21, 22, 23, 24, 25, 13, 26]. For instance, Zhou et al. [20] trained a GAN-based VQAE to learn high-quality (HQ) face priors for blind face restoration. Esser et al. [21] proposed an autoregressive transformer model for high-resolution image generation. MSMC-TTS [22] developed a multi-stage, multi-codebook representation for an advanced text-to-speech system.
Our study regards VQAE as a mechanism for learning and encoding high-frequency facial details, similar to VQ-based approaches. Unlike methods[23, 24, 25, 13, 26] that require additional networks, VQAE inherently possesses noise robustness (see Sec. 3), enabling the recovery of detailed textures from low-quality images. This simplicity makes our method both effective and efficient.
3 Latent Denoising Modeling
In this section, we prove the existence of noise robustness of VQAE via the Lipschitz Continuity theory, where the definition of lipschitz continuity is following:
@@ 3 Latent Denoising Modeling
As depicted in Fig. 1(b), a VQAE is encapsulated by the quadruple $\{\epsilon,\delta,{\bf C},g_{{\bf C}}\}$. Here, $\epsilon$ represents the CNN encoder that maps the image domain $\mathbb{R}^{3\times h\times w}$ to the latent space $\mathbb{R}^{c\times h_{o}\times w_{o}}$. The term $\delta$ signifies the CNN decoder, which reconstructs the image from the latent representation in $\mathbb{R}^{c\times h_{o}\times w_{o}}$. The codebook ${\bf C}=\{{\bf c}_{n}\}_{n=1}^{N}\in\mathbb{R}^{c\times N}$ comprises $N$ anchor vectors. Furthermore, $g_{{\bf C}}$ refers to the channel-wise 1-Nearest Neighbor (1-NN) feature matching operation, responsible for the substitution of latent features with their closest anchor vectors within the codebook. As demonstrated in recent studies [9, 8], the following conclusion can be drawn:
@@
Given a trained VQAE $\{\epsilon,\delta,{\bf C},g_{{\bf C}}\}$, the $\epsilon:~{}\mathbb{R}^{3\times h\times w}\to\mathbb{R}^{c\times h_{o}\times w_{o}}$ is a map with lipschitz continuity such that
@@
We provide the proof in Appendix D. Let ${\bf V}_{up}\subset\mathbb{R}^{T\times 3\times h\times w}$ denote the upscaled version of ${\bf V}_{low}$, where ${\bf V}_{low}$ is the output of Wav2Lip. It is evident that each image in ${\bf V}_{up}$ can be viewed as a degraded HQ image. Given that the primary defect in ${\bf V}_{up}$ is blurring, we naturally assume that this degradation can be modeled as Gaussian blur. Therefore, for any ${\bf I}_{up}\in{\bf V}_{up}$ and ${\bf I}_{high}\in{\bf V}_{high}$, we assume that ${\bf I}_{high}+{\bf\mathcal{N}}={\bf I}_{up}$ always holds, where ${\bf\mathcal{N}}=\text{Gaussian\_blurring}({\bf I}_{high})-{\bf I}_{high}$ represents the domain gap between HQ images and their blurred counterparts. Consequently, by Theorem 3.2, we derive the following results:
$\epsilon({\bf I}_{up})\in\{x\in\mathbb{R}^{c\times h_{o}\times w_{o}}:\|x-\epsilon({\bf I}_{high})\|_{F}\leq L_{\epsilon}\|\mathcal{N}\|_{F}\}.$ (4)
This implies that the $\epsilon(\mathbf{I}_{\text{up}})$ resides within the interior of a hypersphere, with the point $\epsilon(\mathbf{I}_{\text{high}})$ as its center and a radius of $L_{\epsilon}\|\mathcal{N}\|_{F}$, in the measure of the Frobenius norm. Furthermore, for each channel of the latent vector, denoted as $\epsilon(\mathbf{I}_{\text{up}})_{i,j}$ and $\epsilon(\mathbf{I}_{\text{high}})_{i,j}\in\mathbb{R}^{c}$, we have that:
$\|\epsilon({\bf I}_{up})_{i,j}-\epsilon({\bf I}_{high})_{i,j}\|_{2}<L_{\epsilon}\|\mathcal{N}\|_{F}.$ (5)
@@
In a trained VQAE $\{\epsilon,\delta,{\bf C},g_{{\bf C}}\}$, we define $d_{C}={\rm min}\{\|{\bf c}_{i}-{\bf c}_{j}\|_{2}~{}:~{}{\bf c}_{i},{\bf c}_{j}\in{\bf C}~{}and~{}i\neq j\}$ as the minimal distance between anchors of codebook ${\bf C}$, $\gamma$ is the maximal distance of training image latent to closest anchor in all channel, and $L_{\epsilon}$ is the Lipschitz constant of $\epsilon$ when $\|\mathcal{N}\|_{F}<\frac{d_{C}-2\gamma}{2L_{\epsilon}}$ holds, we have:
$g_{\bf C}(\epsilon({\bf I}_{up}))=g_{\bf C}(\epsilon({\bf I}_{high})).$ (6)
@@
This is what we refer to as the latent denoising mechanism of the VQAE. Specifically, the anchors $\{\mathbf{c}_{n}\}_{n=1}^{N}$ partition the latent domain into $N$ high-dimensional Voronoi regions as defined by Equation (1). Let the nearest anchor to $\epsilon(\mathbf{I}_{\text{high}})_{i,j}$ be $\mathbf{c}_{n}$, and $V_{n}$ be its corresponding Voronoi region.  Theorem 3.2 ensures that $\epsilon(\mathbf{I}_{\text{up}})_{i,j}$ remains within $V_{n}$, provided that $\|\mathcal{N}\|_{F}<\frac{d_{C}-2\gamma}{2L_{\epsilon}}$ is satisfied, thereby enabling the $g_{\mathbf{C}}$ operation to correctly assign $\mathbf{c}_{n}$ to $\epsilon(\mathbf{I}_{\text{up}})_{i,j}$. This condition is valid for each channel of $\epsilon(\mathbf{I}_{\text{up}})$, and thus  Theorem 3.3 applies, allowing the decoder $\delta$ to generate HQ faces without loss. The complete proof of  Theorem 3.3 is detailed in  Appendix E.
Figure 2: Display of the noise robustness performance in Macron subject. Please zoom in for better visualization. From top to bottom, there are Gaussian blurred video, denoised video by SOVQAE, ground truth video, denoised video and Gaussian noised video. For each video, we pick 12 continuous frames clip to demonstrate the temporal-consistency of our method.
@@
In Figure 2, we evaluate the denoising performance on a brief talking sequence from the Macron subject, comprising 12 frames. We assess two common video degradation types: Gaussian blurring and Gaussian noise. For each frame, a $192\times 192$ pixel area is randomly selected and subjected to these noise conditions. The increase in PSNR from 30.26 to 35.78 following Gaussian blurring confirms the temporal validity of Theorem 3.3 and supports our assumptions regarding noise $\mathcal{N}$. Even under more semantically destructive Gaussian noise, our method adeptly recovers detail and maintains temporal consistency, with PSNR rising from 16.08 to 30.36. These observations underscore the robustness of SOVQAE to latent noise and affirm its effectiveness.
The aforementioned theoretical analysis substantiates the existence of a denoising mechanism within the VQAE. In our experiments, we observe that a mere 5 minutes of training data suffices for achieving generalization. Furthermore,  Theorem 3.3 posits that the noise tolerance capability is positively correlated with the ratio $\frac{d_{C}-2\gamma}{2L_{\epsilon}}$. To augment this ratio without compromising the denoising efficacy, we introduce an effective regularization loss, as delineated in Equation (11).

### v3

# Lipschitz-Driven Noise Robustness in VQ-AE for High-Frequency Texture Repair in ID-Specific Talking Heads

@@ Abstract
Audio-driven IDentity-specific Talking Head Generation (ID-specific THG) has shown increasing promise for applications in filmmaking and virtual reality. Existing approaches are generally constructed as end-to-end paradigms, and have achieved significant progress. However, they often struggle to capture high-frequency textures due to limited model capacity. To address these limitations, we adopt a simple yet efficient post-processing framework—unlike previous studies that focus solely on end-to-end training—guided by our theoretical insights. Specifically, leveraging the Lipschitz Continuity Theory of neural networks, we prove a crucial noise tolerance property for the Vector Quantized AutoEncoder (VQ-AE), and establish the existence of a Noise Robustness Upper Bound (NRoUB). This insight reveals that we can efficiently obtain an identity-specific denoiser by training an identity-specific neural discrete representation, without requiring an extra network. Based on this theoretical foundation, we propose a plug-and-play Space-Optimized VQ-AE (SOVQAE) with enhanced NRoUB to achieve temporally-consistent denoising. For practical deployment, we further introduce a cascade pipeline combining a pretrained Wav2Lip model with SOVQAE to perform ID-specific THG. Our experiments demonstrate that this pipeline achieves state-of-the-art performance in video quality and robustness for out-of-distribution lip synchronization, surpassing existing identity-specific THG methods. In addition, the pipeline requires only a couple of consumer GPU hours and runs in real time, which is both efficient and practical for industry applications.
@@ 1 Introduction
The generation of photo-realistic, speech-driven, IDentity-specific Talking Head Generation (ID-specific THG) holds significant potential across diverse domains, including filmmaking kim2018deep , virtual reality morishima1998real , and digital avatar creation thies2020neural . The goal of this task is to synthesize ID-specific talking head videos that achieve precise lip synchronization while preserving fine-grained details such as hair, facial wrinkles, moles, eyelashes, and the intricate contours of the lips. These high-frequency details are critical for enhancing realism in synthesized videos.
Existing NeRF-based approaches peng2023synctalk ; guo2021ad ; li2023efficient ; ye2023geneface ; ye2023geneface++ achieve impressive advances in preserving high-fidelity identities. However, these methods typically rely on multilayer perceptrons, which are inherently limited in capturing high-frequency components due to the well-documented spectral bias phenomenon: this bias, as stated in rahaman2019spectral ; basri2020frequency ; yuce2022structured , arises from the pathological eigenvalue distribution of the neural tangent kernel jacot2018neural ; tancik2020fourier , where most eigenvalues are small, limiting convergence on high-frequency components. On the other hand, blind face restoration methods wang2023dr2 ; zhou2022towards ; zhu2022blind , common-used post-processing techniques, are capable of enhancing high-frequency textures, but often suffer from temporal inconsistencies such as texture flickering. While fine-tuning these models on ID-specific data can reduce flickering, it may degrade generalization and incurs additional inefficiencies.
To address these challenges, we propose a simple yet effective post-processing method for ID-specific THG, grounded in rigorous theoretical insights. By analyzing the forward dynamics of VQ-AE  [^1] We use the term VQ-AE to distinguish it from VQGAN esser2021taming , which combines a VQ-AE with a transformer network and emphasizes image synthesis. Here, we focus on the auto-encoder nature of VQ-AE., we uncover the intrinsic noise robustness of VQ-AE and formally derive its Noise Robustness Upper Bound (NRoUB) under the Lipschitz Continuity framework balan2018lipschitz ; zou2019lipschitz (see Sec. 3). This theoretical result implies that training an ID-specific VQ-AE yields an efficient denoiser, where the denoising process remains lossless as long as the noise perturbation stays below the NRoUB. Building on this insight, we introduce a space regularization loss to further improve the NRoUB, resulting in a variant we call Space-Optimized VQ-AE (SOVQAE), which enables temporally consistent denoising (see Fig. 1).
Figure 1: Display of the noise robustness performance in Macron subject. Please zoom in for better visualization. From top to bottom, there are Gaussian blurred video, denoised video by SOVQAE, ground truth video, denoised video and Gaussian noised video. We pick 12 continuous frames clip to demonstrate the temporal-consistency.
To translate this theoretical foundation into a practical THG pipeline, we design a cascade architecture combining a pretrained Wav2Lip model prajwal2020lip —a lightweight and effective lip synchronization model—with an ID-specific SOVQAE (detailed in Sec. 4). Empirical results show that our pipeline achieves new state-of-the-art performance in both video quality and out-of-distribution lip synchronization accuracy, outperforming existing ID-specific THG methods. We attribute this performance to the robust audio-lip synchronization provided by the pretrained model and the high-frequency texture restoration enabled by SOVQAE. Moreover, with half-precision training and inference, our pipeline runs at 30 FPS on a consumer-grade GPU, demonstrating its efficiency and practicality for industrial deployment. In summary, our main contributions are as follows:
  • Theoretical Contribution: We establish the noise robustness of VQ-AE through the lens of Lipschitz continuity theory and derive a formal upper bound, termed the Noise Robustness Upper Bound (NRoUB).
  • Methodological Contribution: We propose SOVQAE, which leverages the derived NRoUB to enable temporally consistent post-processing. Additionally, we introduce an efficient cascade pipeline for identity-specific talking-head generation, combining a pre-trained Wav2Lip model with SOVQAE. The pipeline is simple to train, requiring only a few consumer GPU hours, and operates in real-time inference.
  • Empirical Contribution: Extensive experiments show that our cascade pipeline achieves state-of-the-art performance in video quality and robustness for out-of-distribution lip synchronization, highlighting its practical utility for real-world deployment.
@@ 2 Related Work
Audio-Driven Talking Heads Generation.
As delineated in prajwal2020lip ; shen2023difftalk ; zhang2023dinet ; wang2023seeing ; zhong2023identity ; wang2023lipformer ; stypulkowski2024diffused ; ma2023dreamtalk , one-shot methods harness a plethora of talking video clips to capture the generalizability across multiple faces, thereby attaining commendable lip-synchronization performance. Despite these advancements, these methods continue to grapple with maintaining identity consistency across video frames. To counter this limitation, ID-specific NeRF-based methods peng2023synctalk ; li2023efficient ; ye2023geneface ; guo2021ad ; ye2023geneface++ ; tang2022real are proposed, focusing exclusively on identity-specific data for training. Although this approach mitigates issues of identity instability, the constrained volume of training data hampers their cross-audio synchronization capabilities. To this end, GeneFace ye2023geneface and SyncTalk peng2023synctalk enhance their audio generalization performance by leveraging pre-trained audio-to-landmarks predictors and audio encoders prajwal2020lip respectively. Although these initiatives successfully bolster the cross-audio generalization of NeRF methods, instances of cross-audio desynchronization persist.
Face Restoration.
Blind Face Restoration wang2023dr2 ; zhou2022towards ; gu2022vqfr ; zhu2022blind ; he2022gcfsr ; yang2024pgdiff ; suin2024diffuse ; wang2021towards ; yang2021gan endeavors to rectify diverse degradations in facial imagery captured in non-ideal conditions, including compression, blur, and noise. Given the unknown nature of the specific degradations, this task is inherently ill-posed. The prevalent approach to address this challenge is to integrate prior knowledge within an encoder-decoder framework: GFPGAN wang2021towards utilizes a pre-trained face GAN’s rich priors for blind face restoration. Yang et al. yang2021gan integrate a GAN into a U-shaped DNN for enhanced face restoration. Codeformer wang2021towards uses a learned codebook to predict codes, reducing restoration uncertainty. Through rigorous theoretical analysis and empirical experiments, we establish that the SOVQAE is an effective plug-and-play model for achieving identity-specific, temporally-consistent face restoration, thereby enabling the generation of high-quality talking faces.
Vector Quantization (VQ) Family.
Initially, VQ emerged as a technique within the Signal Processing community gray1984vector ; makhoul1985vector ; cosman1993using ; farvardin1990study . Mathematically, VQ maps signal spaces $\mathbb{R}^{K}$ onto a finite set of Voronoi regions $\{{V_{n}}\}_{n=1}^{N}$, each associated with an anchor vector $\{{\bf c}_{n}\}_{n=1}^{N}\subset\mathbb{R}^{c}$:
@@
where $\|\cdot\|_{2}$ denotes the Euclidean distance. (Fig. 2(a) represents a Voronoi case in 2D space.) Recently, Oord et al. van2017neural integrated the VQ operation into generative models, sparking a trend of discrete latent representation in the field zhou2022towards ; wang2023lipformer ; esser2021taming ; guo2023msmc ; vali2023interpretable ; sadok2024multimodal ; jiang2023text2performer ; rombach2022high ; kaji2023vq .
Our study regards VQ-AE as a mechanism for learning and encoding high-frequency facial details, similar to VQ-based approaches. Unlike methodsvali2023interpretable ; sadok2024multimodal ; jiang2023text2performer ; rombach2022high ; kaji2023vq that require additional networks, VQ-AE inherently possesses noise robustness (Sec.3), enabling the recovery of detailed textures from low-quality images. This simplicity makes our method both effective and efficient.
3 Noise Robustness of VQ-AE
In this section, we derive the noise robustness of VQ-AE via the Lipschitz Continuity theory, where the definition of lipschitz continuity is following:
@@ 3 Noise Robustness of VQ-AE
Figure 2: (a) Illustrates Voronoi regions in 2D and explains the denoising mechanism in VQ space. Given a mapping point $P$, there exists an open disk $\{{\bf x}\in V_{A}\cup V_{B}\cup V_{C}:\|{\bf x}-P\|_{2}<s\}$, where all points in the disk share the same anchor point $A$. (b) Demonstrates the robustness of VQ-AE under slight input disturbances. When image perturbations occur, the VQ mechanism correctly matches the appropriate codebook vector in the latent space, provided the affected vector remains within the correct Voronoi region (e.g., the yellow-red vector in the figure).
As depicted in Fig.2(b), a VQ-AE is encapsulated by the quadruple $\{\epsilon,\delta,{\bf C},g_{{\bf C}}\}$. Here, $\epsilon$ represents the CNN encoder that maps the image domain $\mathbb{R}^{3\times h\times w}$ to the latent space $\mathbb{R}^{c\times h_{o}\times w_{o}}$. The term $\delta$ signifies the CNN decoder, which reconstructs the image from the latent representation in $\mathbb{R}^{c\times h_{o}\times w_{o}}$. The codebook ${\bf C}=\{{\bf c}_{n}\}_{n=1}^{N}\in\mathbb{R}^{c\times N}$ comprises $N$ anchor vectors. Furthermore, $g_{{\bf C}}$ refers to the channel-wise 1-Nearest Neighbor (1-NN) feature matching operation, responsible for the substitution of latent features with their closest anchor vectors within the codebook. As demonstrated in recent studies balan2018lipschitz ; zou2019lipschitz , the following conclusion can be drawn:
@@
Given a trained VQ-AE $\{\epsilon,\delta,{\bf C},g_{{\bf C}}\}$, the $\epsilon:~{}\mathbb{R}^{3\times h\times w}\to\mathbb{R}^{c\times h_{o}\times w_{o}}$ is a map with lipschitz continuity such that
@@
We provide the proof in Appendix.D. Let ${\bf V}_{low}\subset\mathbb{R}^{T\times 3\times h\times w}$ denote the noised version of ${\bf V}_{high}$, where ${\bf V}_{high}$ training video of VQ-AE. Without loss of generality, for any ${\bf I}_{low}\in{\bf V}_{low}$ and ${\bf I}_{high}\in{\bf V}_{high}$, we assume that ${\bf I}_{high}+{\bf\mathcal{N}}={\bf I}_{low}$ holds, where ${\bf\mathcal{N}}$ represents the numerical disturbance caused by any degradation. Consequently, by Theorem 3.2, we derive the following results:
$\epsilon({\bf I}_{low})\in\{x\in\mathbb{R}^{c\times h_{o}\times w_{o}}:\|x-\epsilon({\bf I}_{high})\|_{F}\leq L_{\epsilon}\|\mathcal{N}\|_{F}\}.$ (4)
This implies that the $\epsilon(\mathbf{I}_{\text{low}})$ resides within the interior of a hypersphere, with the point $\epsilon(\mathbf{I}_{\text{high}})$ as its center and a radius of $L_{\epsilon}\|\mathcal{N}\|_{F}$, in the measure of the Frobenius norm. Furthermore, for each channel of the latent vector, denoted as $\epsilon(\mathbf{I}_{\text{low}})_{i,j}$ and $\epsilon(\mathbf{I}_{\text{high}})_{i,j}\in\mathbb{R}^{c}$, we have that:
$\|\epsilon({\bf I}_{low})_{i,j}-\epsilon({\bf I}_{high})_{i,j}\|_{2}<L_{\epsilon}\|\mathcal{N}\|_{F}.$ (5)
@@
In a trained VQ-AE $\{\epsilon,\delta,{\bf C},g_{{\bf C}}\}$, we define $d_{C}={\rm min}\{\|{\bf c}_{i}-{\bf c}_{j}\|_{2}~{}:~{}{\bf c}_{i},{\bf c}_{j}\in{\bf C}~{}and~{}i\neq j\}$ as the minimal distance between anchors of codebook ${\bf C}$, $\gamma$ is the maximal distance of training image latent to closest anchor in all channel, and $L_{\epsilon}$ is the Lipschitz constant of $\epsilon$ when $\|\mathcal{N}\|_{F}<\frac{d_{C}-2\gamma}{2L_{\epsilon}}$ holds, we have:
$g_{\bf C}(\epsilon({\bf I}_{low}))=g_{\bf C}(\epsilon({\bf I}_{high})).$ (6)
@@
This is what we refer to as the noise robustness/tolerance mechanism of the VQ-AE, and we term the $\frac{d_{C}-2\gamma}{2L_{\epsilon}}$ as Noise Robustness Upper Bound (NRoUB). Specifically, the anchors $\{\mathbf{c}_{n}\}_{n=1}^{N}$ partition the latent domain into $N$ high-dimensional Voronoi regions as defined by Eq. (1). Let the nearest anchor to $\epsilon(\mathbf{I}_{\text{high}})_{i,j}$ be $\mathbf{c}_{n}$, and $V_{n}$ be its corresponding Voronoi region. Theorem 3.2 ensures that $\epsilon(\mathbf{I}_{\text{low}})_{i,j}$ remains within $V_{n}$, provided that $\|\mathcal{N}\|_{F}<\frac{d_{C}-2\gamma}{2L_{\epsilon}}$ is satisfied, thereby enabling the $g_{\mathbf{C}}$ operation to correctly assign $\mathbf{c}_{n}$ to $\epsilon(\mathbf{I}_{\text{low}})_{i,j}$. This condition is valid for each channel of $\epsilon(\mathbf{I}_{\text{low}})$, and thus Theorem 3.3 applies, allowing the decoder $\delta$ to generate high quality faces without loss. The complete proof of Theorem 3.3 is detailed in Appendix E.
@@
As shown in Fig. 1, we validate the noise robustness on a brief talking sequence with two common video degradation types: Gaussian blurring and Gaussian noise. For each frame, a $192\times 192$ pixel area is randomly selected and subjected to these noise conditions. The increase in PSNR from 30.26 to 35.78 following Gaussian blurring confirms the temporal validity of Theorem 3.3. Even under more semantically destructive Gaussian noise, our method adeptly recovers detail and maintains temporal consistency, with PSNR rising from 16.08 to 30.36. These observations underscore the robustness of VQ-AE to noise and affirm its effectiveness.
Application insights for High-Frequency Texture Repair in ID-specific THG.
On the one hand, in ID-specific THG, generating realistic audio-driven videos with high-frequency details is challenging to existing end-to-end methods. A common way to enhance face details is doing face restoration. However existing face restoration models fall short in temporal-consistency. Hence, a feasible post-processing method for ID-specific THG is to enhance texture with temporal-consistency.
On the other hand, our aforementioned theoretical analysis highlights two important and interesting conclusions:
  • An ID-specific denoiser can be obtained through training the ID-aware neural discrete representation (VQ-AE);
  • The denoising performance can be temporal-consistent, when the noise disturbance is below the NRoUB.
Therefore, if we can increase the NRoUB of VQ-AE, we can obtain the ideal post-processing model for ID-specific THG. To this end, we introduce a space regularization loss to enhance the NRoUB of VQ-AE (termed as SOVQAE) in next section, and illustrate its combination with pretrained Wav2Lip showing its application potential.
4 Space-Optimized VQ-AE and Application
In this section, we firstly introduce a feasible space regularization loss to enhance the NRoUB, building our SOVQAE, and illustrate how the pretrained Wav2Lip combines with our SOVQAE to build state-of-the-art ID-specific THG pipeline.